OpenStudy (pandalife):

MEDAL + FAN a circle has its center at (-3,2) and a radius of 5 units.what is the equation of the CIRCLE?

OpenStudy (bibby):

what is the generall form equation of a circle?

OpenStudy (pandalife):

THATS all they ask me

OpenStudy (bibby):

whoops, meant center-radius form

OpenStudy (bibby):

\(\large (x-h)^2+(y-k)^2=r^2\) where r is the radius and the center is (h,k)

OpenStudy (pandalife):

can we do it together?

OpenStudy (bibby):

you haven't even bought me dinner yet.

OpenStudy (bibby):

just plug in the information they've given you. you have the center (h,k) = (-3,2) plug that into the \(\large (x-h)^2+(y-k)^2=r^2\) equation

OpenStudy (bibby):

for example if our circle was at 1,3 it'd look like \((x-1)^2 + (y-3)^2\) (2,-4) would look like \((x-2)^2 + (y--4)^2=(x-2)^2 + (y+4)^2\)

OpenStudy (pandalife):

(x--3)2 + (y-2)2=?

OpenStudy (bibby):

yeah, good job. now what is r?

OpenStudy (pandalife):

5

OpenStudy (pandalife):

??

OpenStudy (bibby):

yeah!!

OpenStudy (pandalife):

so what do we do now?

OpenStudy (bibby):

nothing. we have the equation

OpenStudy (pandalife):

can you write it out for me so i can see how it looked squared

OpenStudy (bibby):

just in case it isn't clear, \((x--3)^2 + (y-2)^2=5^2\) simplify it yourself

OpenStudy (pandalife):

no help and i paid for dinner?

OpenStudy (bibby):

you bought dinner just to get into my reputably loose pants

OpenStudy (bibby):

what is (x--3)? what is 5^2?

OpenStudy (pandalife):

x3+25?

OpenStudy (bibby):

x3?

OpenStudy (pandalife):

?

OpenStudy (bibby):

x3+25? what does that even mean? I asked you for 2 different expressions not a sum

OpenStudy (pandalife):

can you help me

OpenStudy (bibby):

I've given you the answer bro \(\huge (x--3)^2 + (y-2)^2=5^2\)

OpenStudy (pandalife):

thanks